# Conditional regression for the Nonlinear Single-Variable Model

### v1

@@ 2.4 The main algorithm and guarantees on the estimator it produces
    Algorithm 1 summarizes the construction of the proposed estimator of the regression function $F$. We prove that the mean squared error of our estimator is the sum of the estimation error and the curve approximation error. The estimation error decays the one-dimensional min-max optimal nonparametric learning rate, up to a saturation level dependent on quantities expected to be small, or even zero, as we discuss in detail below.
@@
    Assume that (X subG), (Y subG), ($\bm{\zeta}$ subG), ($\bm{\gamma_{1}}$), (LCV), ($\bm{\omega_{f}}$), and (P) hold true, and that $f\in\mathcal{C}^{s}(\mathbb{R}^{1})$ with $s\in\left[\frac{1}{2},2\right]$. Let $C_{\gamma,f},M^{*},l_{\max}$ be the constants specified in (4) below, and $C$ a universal constant. Then, for $n$ such that $\frac{n}{\log^{{3/2}}n}\gtrsim C_{\gamma,f}\frac{C_{f}\mathrm{len}_{\gamma}}{C_{f}^{\prime}\sigma_{\gamma}}$, if we choose $l^{*},j^{*}$ as
@@
    $\mathbb{E}\left[\left|{\widehat{F}}(X)-F(X)\right|^{2}\right]\lesssim C_{1}(f,\gamma,\rho_{X},\sigma_{\zeta},d)n^{-\frac{2s}{2s+1}}{\log n}+C_{2}(f,\gamma,\rho_{X},\sigma_{\zeta},d)$
@@
    The above theorem states that our estimator achieves the min-max optimal rate for the one-dimensional nonparametric regression, with an additional approximation error of magnitude $\mathcal{O}(\sigma_{\zeta}^{2})$ for functions that are both Lipschitz and monotone, therefore defeating the curse of dimensionality by exploiting the compositional structure of $F$, even if the inner function is nonlinear (unlike the single- and multi-index model) and not particularly smooth (and its regularity does not scale with the ambient dimension). It is worth mentioning that the result of Theorem 3 is scaling invariant in $X$ and $Y$.
@@
      the minimal number of samples $n$ required is only a low-order polynomial in $d$, $\mathrm{len}_{\gamma}$, and $\left\|\gamma^{\prime\prime}\right\|$. This is not exponential in ambient dimension $d$ (unless in the extreme case where $\gamma$ has length or curvature growing exponentially in $d$, e.g. a space-filling curve).
@@
      the assumption that $f$ is coarsely monotone. While similar conditions have appeared in other works using conditional regression (see Lanteri et al., 2022, for a discussion), and even in recent approaches based on gradient descent for the single- and multi-index models (see e.g., Damian et al., 2024; Arous et al., 2021; Bietti et al., 2023, and references therein) the “amount of oscillation” of $f$ appears prominently and imposes additional sampling requirements (e.g., scaling as $d^{\mathcal{O}(L)}$ where $L$ is the number of vanishing moments in the Hermite polynomial basis, albeit it is not clear to us that such assumptions and sampling requirements are sharp);
@@
      the second shortcoming is that we have a second additive term, which we call curve approximation error, which does not vanish as $n$ tends to infinity. It is typically small, e.g., if external noise $\sigma_{\zeta}$ or the curvature $\left\|\gamma^{\prime\prime}\right\|$ is small, and in fact it will vanish when external noise $\sigma_{\zeta}$ or curvature $\left\|\gamma^{\prime\prime}\right\|$ vanishes. In the limiting case where the underlying curve $\gamma$ is a straight line segment, we have no curve approximation error, recovering the results of (Lanteri et al., 2022). Otherwise, the nonzero approximation error exists because we are using first-order approximations in the estimation of the local directions of the curve, and with noisy observations our technique for constructing the estimator does not allow us to consider local pieces below the scale of the noise. While it seems possible to use approximations of higher order, they would require significant additional computational complexity and more refined statistical analysis, possibly with minimal impact in practical applications, and are left to future work.
@@
    Assume that (X subG), (Y subG), ($\bm{\zeta}$ subG), ($\bm{\gamma_{1}}$), (LCV), ($\bm{\omega_{f}}$), and (P) hold true. Assume that there is no observational noise, i.e., $\zeta\equiv 0$ almost surely, that the link function $f$ is perfectly monotone, i.e., $\omega_{f}=0$, and that $f\in\mathcal{C}^{s}$ with $s\in\left[\frac{1}{2},2\right]$. With $C_{\gamma,f}$ in (4), when $\frac{n}{\log^{3/2}n}\gtrsim C_{\gamma,f}\frac{C_{f}\mathrm{len}_{\gamma}}{C_{f}^{\prime}\sigma_{\gamma}}$, if we choose $l^{*}=\frac{1}{C_{\gamma,f}}\frac{n}{\log^{3/2}n}$ and $j^{*}=C$, then the estimation error of Algorithm 1 satisfies
@@
    The better rate $\mathcal{O}((n^{-2}\log^{3}n)^{(s\wedge 1)})$ is a consequence of having zero observational noise. Bauer et al. (2017) proves that, in the case of the $L^{\infty}$-norm, the min-max rate of nonparametric regression for functions in $\mathcal{C}^{s}(\mathbb{R}^{d})$ with noiseless observations is $(\log n/n)^{s/d}$. The rate on the mean squared error in Theorem 4 agrees with this when $s\leqslant 1$ and $d=1$, as if $\gamma$ (and therefore $\Pi_{\gamma}$) were known. Our rate is suboptimal for $s>1$, as we only perform a first-order approximation of the underlying curve and do not estimate higher-order parameters such as the curvature. Our estimator avoids the curse of dimensionality by exploiting the compositional structure of $F$, even if the inner function is (unknown and) nonlinear (unlike in the single- and multi-index model) and not particularly smooth (and its regularity does not scale with the ambient dimension). A key difference between Theorem 3 and Theorem 4 is that there is no curve approximation error in the latter: noiseless observations allow us to perform an unlimited amount of partitioning, obtaining “thin” slices, provided that enough samples are available (in order to control the variance of the objects we estimate in each slice).
@@
    Assume that (X subG), (Y subG), ($\bm{\zeta}$ subG), ($\bm{\gamma_{1}}$), ($\bm{\omega_{f}}$), and (SC) hold true. When $\frac{n}{\log^{1/2}n}\gtrsim\frac{C_{Y}R_{0}^{7}d^{3}}{C_{f}^{6}\max(\sigma_{\zeta},\omega_{f})^{7}}$, if we choose $l^{*}=\frac{C_{Y}R_{0}}{\max(\sigma_{\zeta},\omega_{f})}$ and $j^{*}=C$, the estimation error of Algorithm 1 satisfies
@@ 3 Analysis of the Estimator
    1. (X subG)
      $X$ has sub-Gaussian distribution with variance proxy $R_{0}^{2}$, and has a density function $\rho_{X}$ with $\left[\rho_{X}\right]_{\mathcal{C}^{2}}<\infty$.
    1. (Y subG)
@@
    Assumption ($\bm{\gamma_{1}}$) implies that for any $t_{0}\in[0,\mathrm{len}_{\gamma}]$, the conditional mean is on the curve
@@
    $=\mathbb{E}\left[(\gamma(t)-\mu)(\gamma(t)-\mu)^{\intercal}\mid t\in T\right]+\mathbb{E}[\sigma_{\gamma}(t)^{2}\mid t\in T]\mathbb{I}_{d-1}-\mathbb{E}[\sigma_{\gamma}(t)^{2}\gamma^{\prime}(t)\gamma^{\prime}(t)^{\intercal}\mid t\in T]$
@@
    1. (P)
      The density $\rho_{X}$ has compact support. Let $\rho_{v}$ denote the push-forward measure of the distribution of $X$ along the map $x\mapsto\langle x,v\rangle$. Then there exists constant $c>0$ such that, for every $v\in\mathbb{S}^{d-1}$ and for every interval $I\subseteq\operatorname{\mathrm{supp}}\rho_{v}$, we have $\rho_{v}(I)\geqslant c|I|$ and $\mathrm{Var}[\langle v,X\rangle|\langle v,X\rangle\in I]\geqslant c|I|^{2}$.
    Assumption ($\bm{\gamma_{1}}$) implies that for $t_{0}\in[0,\mathrm{len}_{\gamma}]$, $\sigma_{\gamma}(t_{0})<\min(R_{0},\mathrm{reach}_{\gamma}/\sqrt{d})$; (LCV) implies that $\sigma_{\gamma}\geqslant 2C_{f}\sigma_{\zeta}$; ($\bm{\omega_{f}}$) implies that ${C_{f}^{\prime}}C_{Y}R_{0}\leqslant\mathrm{len}_{\gamma}\leqslant C_{f}C_{Y}R_{0}$.
@@ 3.1 Estimation of slice parameters with Assumption (LCV)
    ${\mu_{l,h}^{b}}:=\mathbb{E}\left[X|Y\in{R_{l,h}},\mathcal{B}\right],\ \ \ {\Sigma_{l,h}^{b}}:=\operatorname{{Cov}}\left[X|Y\in{R_{l,h}},\mathcal{B}\right]$
@@
    $\mathcal{B}_{i}:=\{\left\|X_{i}\right\|\leqslant C_{X}\sqrt{d}R_{0},|Y_{i}|\leqslant C_{Y}R_{0}\}$
@@
    The random variable $n^{b}$, assuming (X subG) and (Y subG), is larger than a constant fraction of $n$ with high probability (see Lemma 21). The sample counterparts of ${\mu_{l,h}^{b}}$ and ${\Sigma_{l,h}^{b}}$ are
@@
    Suppose ($\bm{\zeta}$ subG) and ($\bm{\omega_{f}}$) hold true. Let $T\subseteq R$ be a bounded interval with $|T|\geqslant\max(\sigma_{\zeta},\omega_{f})$. Then:
@@
    Suppose ($\bm{\zeta}$ subG), ($\bm{\gamma_{1}}$) and ($\bm{\omega_{f}}$) hold true. Let $T\subseteq R$ be a bounded interval with $|T|\geqslant\max(\sigma_{\zeta},\omega_{f})$. Then:
@@
    Suppose (X subG), (Y subG), ($\bm{\zeta}$ subG), ($\bm{\gamma_{1}}$), (LCV), and ($\bm{\omega_{f}}$) hold true. Then, for every $l$ such that $l\gtrsim C_{f}C_{Y}R_{0}/\sigma_{\gamma}$, and $|{R_{l,h}^{b}}|\geqslant\max(\sigma_{\zeta},\omega_{f}),h\in\mathcal{H}_{l}^{b}$, we have
@@
    Suppose (X subG), (Y subG), ($\bm{\zeta}$ subG), ($\bm{\gamma_{1}}$), (LCV), and ($\bm{\omega_{f}}$) hold true. Then, for every $l$ such that $l\gtrsim C_{f}C_{Y}R_{0}/\sigma_{\gamma}$, and $|{R_{l,h}^{b}}|\geqslant\max(\sigma_{\zeta},\omega_{f})$ for all $h\in\mathcal{H}_{l}^{b}$, for every $\epsilon>0$ and $\tau\geqslant 1$, if $n$ is sufficiently large so that $\frac{n}{\sqrt{\tau\log n}}\gtrsim\left(\frac{C_{f}C_{Y}R_{0}}{\sigma_{\gamma}}\right)^{2}dl(t+\log d+\log l)$, we have
    $\mathbb{P}\left\{\exists h\in{\mathcal{H}_{l}^{b}}:\left|\langle{v_{l,h}^{b}},{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\rangle\right|\gtrsim C_{Y}C_{f}R_{0}\sqrt{t+\log l+\log d}\sqrt{\frac{\sqrt{\tau\log n}}{nl}}\ \bigg{|}\ {n_{l,h}^{b}}\right\}\lesssim e^{-t}+n^{-\tau}$
    $\mathbb{P}\left\{\exists h\in{\mathcal{H}_{l}^{b}}:\left\|{{\widehat{v}_{l,h}^{b}}}-{{v_{l,h}^{b}}}\right\|\gtrsim C_{Y}C_{f}R_{0}^{2}\sigma_{\gamma}^{-2}\sqrt{t+\log l+\log d}\sqrt{\frac{d\sqrt{\tau\log n}}{nl}}\ \bigg{|}\ {n_{l,h}^{b}}\right\}\lesssim e^{-t}+n^{-\tau}$
@@
    $\mathbb{E}\left[\left\|{\widehat{v}_{l,h}^{b}}-{v_{l,h}^{b}}\right\|^{2p}\right]\lesssim C(p)(C_{Y}C_{f}R_{0}^{2}\sigma_{\gamma}^{-2})^{2p}(d\log d)^{p}\left(\frac{\log n\log l}{nl}\right)^{p}$

### v2

@@ 2.4 The main algorithm and guarantees on the estimator it produces
    Algorithm 1 summarizes the construction of the proposed estimator of the regression function $F$. We prove, under assumptions that are detailed in Section 3, that the mean squared error of our estimator is the sum of the estimation error and the curve approximation error. The estimation error decays at the one-dimensional min-max optimal nonparametric learning rate, up to log factors and, more importantly, up to a saturation level dependent on quantities expected to be small, or even zero, as we discuss in detail below.
@@
[…]
@@
    $\mathbb{E}\left[\left|{\widehat{F}}(X)-F(X)\right|^{2}\right]\lesssim C_{1}(f,\gamma,\rho_{X},\sigma_{\zeta},d)n^{-\frac{2s}{2s+1}}{\log n}+C_{2}(f,\gamma,\rho_{X},\sigma_{\zeta},d)\ \,.$
@@
    The above theorem states that our estimator achieves the min-max optimal rate, (up to log factors), for the one-dimensional nonparametric regression, with an additional approximation error of magnitude $\mathcal{O}(\sigma_{\zeta}^{2})$ for functions that are both Lipschitz and monotone, therefore defeating the curse of dimensionality by exploiting the compositional structure of $F$, even if the inner function is nonlinear (unlike the single- and multi-index model) and not particularly smooth (and its regularity does not scale with the ambient dimension). It is worth mentioning that the result of Theorem 3 is scaling invariant in $X$ and $Y$.
@@
      the minimal number of samples $n$ required is only a low-order polynomial in $d$, $\mathrm{len}_{\gamma}$, and $\left\|\gamma^{\prime\prime}\right\|$. This is not exponential in ambient dimension $d$ (unless in the extreme case where $\gamma$ has length or curvature growing exponentially in $d$, e.g., a space-filling curve);
@@
      the regularity assumptions on both $f$ and $\gamma$ are independent of the dimension $d$;
    5. (v)
@@
      the assumption that $f$ is coarsely monotone. While similar conditions have appeared in other works using conditional regression (see Lanteri et al., 2022, for a discussion), and even in recent approaches based on gradient descent for the single- and multi-index models (see e.g., Damian et al., 2024; Arous et al., 2021; Bietti et al., 2023, and references therein) the “amount of oscillation” of $f$ appears prominently and imposes additional sampling requirements (e.g., scaling as $d^{\mathcal{O}(L)}$ where $L$ is the number of vanishing moments in the Hermite polynomial basis, albeit it is not clear to us that such assumptions and sampling requirements are sharp). At this point we do not know if this limitation can be overcome by estimators similar to ours with a cost still polynomial in $d$. In similar and related models, albeit not applicable to the nonlinear variable model introduced here, the estimators in Bach (2017) and Chen and Meka (2020) do not require restrictive assumptions, so it may indeed the case that this restrictions are due to the specific family of estimators we introduce here.
@@
      the second shortcoming is that we have a second additive term, which we call curve approximation error, which does not vanish as $n$ tends to infinity. It is typically small, e.g., if external noise $\sigma_{\zeta}$ or the curvature $\left\|\gamma^{\prime\prime}\right\|$ is small, and in fact it will vanish when external noise $\sigma_{\zeta}$ or curvature $\left\|\gamma^{\prime\prime}\right\|$ vanishes. In the limiting case where the underlying curve $\gamma$ is a straight line segment, we have no curve approximation error, recovering the results of (Lanteri et al., 2022). Otherwise, the nonzero approximation error appears here because we are using first-order approximations in the estimation of the local directions of the curve, and with noisy observations our technique for constructing the estimator does not allow us to consider local pieces below the scale of the noise, no matter what the sample size. While it seems possible to use approximations of higher order, increasing with $n$, would allow us to obtain a term vanishing in $n$, they would require significant additional computational complexity and more refined statistical analysis, possibly with minimal impact in practical applications, and are left to future work. At this point we do not know if this is an artifact of the use of inverse-regression techniques, or if any polynomial (in $n$ and $D$) time algorithm would necessarily require similar hypotheses.
@@
[…]
@@
    The better rate $\mathcal{O}((n^{-2}\log^{3}n)^{(s\wedge 1)})$ is a consequence of having zero observational noise. Bauer et al. (2017) proves that, in the case of the $L^{\infty}$-norm, the min-max rate of nonparametric regression for functions in $\mathcal{C}^{s}(\mathbb{R}^{d})$ with noiseless observations is $(\log n/n)^{s/d}$. The rate on the mean squared error in Theorem 4 is consistent with this rate in $L^{\infty}$ when $s\leqslant 1$ and $d=1$, as if $\gamma$ (and therefore $\Pi_{\gamma}$) were known. Our rate is suboptimal for $s>1$, as we only perform a first-order approximation of the underlying curve and do not estimate higher-order parameters such as the curvature. Our estimator avoids the curse of dimensionality by exploiting the compositional structure of $F$, even if the inner function is (unknown and) nonlinear (unlike in the single- and multi-index model) and not particularly smooth (and its regularity does not scale with the ambient dimension). A key difference between Theorem 3 and Theorem 4 is that there is no curve approximation error in the latter: noiseless observations allow us to perform an unlimited amount of partitioning, obtaining “thin” slices, provided that enough samples are available (in order to control the variance of the objects we estimate in each slice).
@@
[…]
@@ 3 Analysis of the Estimator
[…]
      $X$ has sub-Gaussian distribution with variance proxy $R_{0}^{2}$, and has a density function $\rho_{X}$ which is $\mathcal{C}^{2}$ with $\left[\rho_{X}\right]_{\mathcal{C}^{2}}<\infty$ and has compact support.
[…]
@@
    This assumption is satisfied, for example, by a natural generative model where a point $\gamma(t)$ is sampled on the curve $\gamma$ and, conditional to that, a point $X$ is produced according to (1) with $Z_{d-1}$ a sub-Gaussian distribution as in ($\bm{\gamma_{1}}$). Note that this does not imply that, overall, the points in the normal directions to the curve are uniformly or isotropically distributed. Assumption ($\bm{\gamma_{1}}$) implies that for any $t_{0}\in[0,\mathrm{len}_{\gamma}]$, the conditional mean is on the curve
@@
    $=\mathbb{E}\left[(\gamma(t)-\mu)(\gamma(t)-\mu)^{\intercal}\mid t\in T\right]+\mathbb{E}[\sigma_{\gamma}(t)^{2}\mid t\in T]\mathbb{I}_{d}-\mathbb{E}[\sigma_{\gamma}(t)^{2}\gamma^{\prime}(t)\gamma^{\prime}(t)^{\intercal}\mid t\in T]\ \,.$
@@
    1. ($\bm{\text{P}_{\gamma}}$)
      Let $\rho_{t}$ denote the push-forward measure of the distribution of $X$ along the map $x\mapsto\Pi_{\gamma}(x)\in[0,\mathrm{len}_{\gamma}]$. Then there exists constant $c>0$ such that for every interval $I\subseteq[0,\mathrm{len}_{\gamma}]$ with $|I|\leqslant\sigma_{\gamma}$, we have $\rho_{t}(I)\geqslant c|I|$ and $\mathrm{Var}[\Pi_{\gamma}(X)|\Pi_{\gamma}(X)\in I]\geqslant c|I|^{2}$.
    Assumption ($\bm{\text{P}_{\gamma}}$) is rather weak: it states that if we project the random variable $X$ onto the underlying curve $\gamma$ then the corresponding distribution on the curve has a nontrivial lower-bound on its density.
    It is worth noticing that the above assumptions put some restriction on the parameters of the curve $\gamma$: Assumption ($\bm{\gamma_{1}}$) implies that for $t_{0}\in[0,\mathrm{len}_{\gamma}]$, $\sigma_{\gamma}(t_{0})<\min(R_{0},\mathrm{reach}_{\gamma}/\sqrt{d})$; (LCV) implies that $\sigma_{\gamma}\geqslant 2C_{f}\sigma_{\zeta}$; ($\bm{\omega_{f}}$) implies that ${C_{f}^{\prime}}C_{Y}R_{0}\leqslant\mathrm{len}_{\gamma}\leqslant C_{f}C_{Y}R_{0}$.
@@ 3.1 Estimation of slice parameters with Assumption (LCV)
    ${\mu_{l,h}^{b}}:=\mathbb{E}\left[X|Y\in{R_{l,h}},\mathcal{B}\right],\ \ \ {\Sigma_{l,h}^{b}}:=\operatorname{{Cov}}\left[X|Y\in{R_{l,h}},\mathcal{B}\right]\ \,.$
@@
    $\mathcal{B}_{i}:=\{\left\|X_{i}\right\|\leqslant C_{X}\sqrt{d}R_{0},|Y_{i}|\leqslant C_{Y}R_{0}\}\ \,,$
@@
[…]
@@
    Suppose ($\bm{\zeta}\in\psi_{2}$) and ($\bm{\omega_{f}}$) hold true. Let $T\subseteq R$ be a bounded interval with $|T|\geqslant\max(\sigma_{\zeta},\omega_{f})$. Then:
@@
    Suppose ($\bm{\zeta}\in\psi_{2}$), ($\bm{\gamma_{1}}$) and ($\bm{\omega_{f}}$) hold true. Let $T\subseteq R$ be a bounded interval with $|T|\geqslant\max(\sigma_{\zeta},\omega_{f})$. Then:
@@
[…]
@@
[…]
    $\mathbb{P}\left\{\exists h\in{\mathcal{H}_{l}^{b}}:\left|\langle{v_{l,h}^{b}},{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\rangle\right|\gtrsim C_{Y}C_{f}R_{0}\sqrt{t+\log l+\log d}\sqrt{\frac{\sqrt{\tau\log n}}{nl}}\ \bigg{|}\ {n_{l,h}^{b}}\right\}\lesssim e^{-t}+n^{-\tau}\ \,;$
    $\mathbb{P}\left\{\exists h\in{\mathcal{H}_{l}^{b}}:\left\|{{\widehat{v}_{l,h}^{b}}}-{{v_{l,h}^{b}}}\right\|\gtrsim C_{Y}C_{f}R_{0}^{2}\sigma_{\gamma}^{-2}\sqrt{t+\log l+\log d}\sqrt{\frac{d\sqrt{\tau\log n}}{nl}}\ \bigg{|}\ {n_{l,h}^{b}}\right\}\lesssim e^{-t}+n^{-\tau}\ \,.$
@@
    $\mathbb{E}\left[\left\|{\widehat{v}_{l,h}^{b}}-{v_{l,h}^{b}}\right\|^{2p}\right]\lesssim C(p)(C_{Y}C_{f}R_{0}^{2}\sigma_{\gamma}^{-2})^{2p}(d\log d)^{p}\left(\frac{\log n\log l}{nl}\right)^{p}\ \,.$